

********************************************************************
UARS WEEKLY STATUS REPORT - 28 JUL 2000
********************************************************************
***********
SOLSTICE
***********
This week's large solar pointing excursions (15 arcminutes at times) are
still within our calibrations so have minimal affect our solar data.
Additionally, we continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder
operations.
SOLSTICE is in good health, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. Level 2 processing is halted due to bugs in the OA_SSPP_ATT
service. CDHF and CSC personnel are working on the problem.
***********
SUSIM
***********
SUSIM successfully conducted its yearly calibration scans involving
all 4 deuterium lamps. Otherwise, the instrument continues to gather
a normal set of daily solar scans scheduled for periods of predicted
telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
HALOE has been off during this full graze period. HALOE is schedule to
turn on beginning of day August 2, 2000.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Processing of instrument Daily Activity Plans continues with no
problems. We are continuing to coordinate with the Flight Operations
Team and to assist the Project Scientist in scheduling of instrument
operations. The schedule for instrument operations during the next
yaw period (11 August to 17 September) has been distributed and posted
on the web.(http://umpgal.gsfc.nasa.gov/)
The MPG attended a meeting to discuss the status of Fixed Head Star Tracker 1
on Wednesday, July 26, 2000.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
2 days SEN (5.0) PROD (1992) (2R-2C=0L)
Data created in the month of May, 2000 has been loaded into the DAAC
Transfer Tables for HALOE, NMC, and UKMO. DAAC has been notified.
UARS CDHF REHOST:
- Intentions are on July 31,2000 to begin running current day processing,
Level 0 and NRT Ingest on both the present production and rehost systems.
Approximate duration 2 weeks. After this period (assuming no problems found)
we would then like PI/SW reps to use the REHOST system the primary system for
LO ,NRT and current day processing. PAC and Production will notify PI software
reps regarding data transfer and account access on the Rehost.
- Successful processing on Rehost System:
Attitude Extraction, Observatory Log, Sensor Extraction, HALOE, HRDI,
MLS, SUSIM, Solstice, and WINDII.
- The HRDI group has validated the V4.4 output product created
on the Rehost System.
- A re-linking of all the PI software tested has been done.
- United Kingdom Meteorological Office is sending daily Correlative
(UKMO) data to the Rehost System. Needed for HALOE and HRDI production
processing. Same data continues to be sent to the current system.
No differences when comparing rehost and current files.
- A daily Auxiliary (SOL) file is now being received on the Rehost System.
Needed for HALOE Production processing. Same data continues to be sent
to the current system. Found no differences when comparing rehost and
current files.
- Production control has tested all PI software that has been re-compiled
and re-linked with the rebuilt UCSS shared image software. Test days
were run against UARS DAYS 3206 - 3213 that were electronically ingested
on the Rehost System. Results are being compared against same test days
that were successfully processed before UCSS rebuild.
***********
FOT
***********
REPORTING PERIOD
This report covers 15 July 2000 (Orbit 48332, GMT Day 197) through 21 July 2000
(Orbit 48436 , GMT Day 203).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 200/2346 48391 Off
WINDII 201/0126 48392 On
PEM 202/2210 48420 On, 7w
PEM 203/2032 48434 On, PMON
WINDII 203/2351 48436 Off
AXIS-1 204/2353 48451 Off
The beta angle ranged between the angles of 38.5 degrees to 67.2 degrees for
this report period. The beta angle is now increasing toward a maximum angle of
+76 degrees on 25 July 2000 (DOY 207).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.7 to 1.0 degrees C. Additional battery performance data for each
battery on 14 July 2000 and 21 June 2000 are provided in Appendix D.
The clock error ranged from -11.7 to 11.6 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 02
% ALERTS = 01
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 09
% WINDII = 03
% PEM = 03
% HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 00
Two (2) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
- Transponder A: 00/145 - 2287.497008
- Transponder B: 00/152 - 2287.498920
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/202 show the observatory data loss to be 296 hours, 38
minutes, 23.737 seconds (an increase of 1 minute 20.896 seconds since the last
report period). This is a 0.40 percent data loss which equals a 99.60 percent
data capture for the mission.
The increase of 1 minute 20.896 seconds occurred over two (2) days:
- 0 minutes 55.728 seconds on Day 189 due a late acquisition,
- 0 minutes 25.168 seconds on Day 193 due to a data hit.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
UARS continues to operate this week using FHST-1 and the FSS as its update
sensors for attitude control. FHST-1 has shown, as expected, degraded
performance. A scale factor calibration was made to FHST-1 on 17 July 2000 and
an alignment calibration on 21 July 2000. These calibrations were given to the
FOT by FDF and should help to correct the error in the star tracker. The FOT
will continue to monitor the data to see how much of an improvement the
calibrations have made to UARSU pointing.
The FOT has delayed the testing of the new TMONs until the FHST-2 anomaly is
resolved. No time frame has been set as to when testing will resume.
The FOT has begun writing the online procedure for safehold should it happen
before the new TMONs are in place or during and OBC halt. This procedure will
be tested on the simulator and put online for immediate execution should an
anomalous condition take place.
MEETINGS
The FOT attended a meeting to discuss FHST-1 on 19 July 2000.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for August 14, 2000 (DOY 227). This
maneuver will be a Forward to Reverse Flight maneuver (Northern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 21 July 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
48379 200/0422 00-348 NBTR PLAYBACK
48380 200/0603 00-348 NBTR NOOPS D200
48391 200/2346 00-351 HALOE Power OFF
48392 201/0126 00-352 WINDII Power ON
48393 201/0309 00-353 WINDII WINDII ALERT CLEAR
48394 201/0424 00-354 NBTR PLAYBACK
48399 201/1147 00-355 OBC Clk Adj fr 0.00 to 20.6 msec/day
48409 202/0426 00-356 NBTR PLAYBACK
48420 202/2210 00-299 PEM "Power ON, 7w"
48424 203/0428 00-357 NBTR PLAYBACK
48433 203/1809 00-358 OBC Clk Adj 20.6 to 41.2 msec/day
48434 203/2032 00-359 PEM Power ON - DAY ONLY
48435 203/2129 00-360 OBC FHST-1 ALIGNMENT
48436 203/2351 00-362 WINDII Power OFF
48439 204/0430 00-361 NBTR PLAYBACK
48451 204/2353 00-363 AXIS-1 Power OFF
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 21 July 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPSEQBIMON Y-HI PSE QB CURRENT MON 48381
WIISUTEMP Y-LO WINDII ISU TEMP 48381
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
WIMCCNOK WINDII SCAN MIRROR IMPROPER FUNCTION 48393
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 21 July 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
48399 201/12:00 (19 July 00) from 0.00 to 20.6 msec/day
48433 203/18:27 (21 July 00) from 20.6 to 41.2 msec/day
___________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 21 July 2000
14 July 2000 - GMT Day 196 Beta = 38.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ---------- ------ ------- ------
1 +56.0/+33.6 -3.05/-4.31 +0.0/+0.0 15.5 0.000 0.0
2 +50.4/-56.0 +5.79/+4.51 +28.0/-15.2 25.0 1.030 22.0
3 +22.4/-33.6 +5.15/+3.23 +27.2/-16.4 25.0 1.030 22.0
21 July 2000 - GMT Day 203 Beta = 67.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ---------- ------ ------- ------
1 +56.0/+39.2 -3.05/-4.62 +0.0/+0.0 15.5 0.000 0.0
2 +16.8/-11.2 +3.55/+1.97 +19.2/-15.2 26.1 1.040 17.4
3 +22.4/-11.2 +2.60/+1.34 +19.2/-15.2 25.9 1.030 17.1
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 21 July 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-026 FOT see list June 00 14170 Generic Late Acq.
00-032 FHST 48211 188/2242 N/A FHST 2 Anomaly
SEVNTFW00/29